OpenStudy (anonymous):

3sinx = 2cos^2x solve.

OpenStudy (anonymous):

Graph them on your calculator my friend and find the intersecting points ;)

OpenStudy (anonymous):

no i have to solve them without a calculator

OpenStudy (anonymous):

Just clarifying that 2cos^2x is\[2\cos^2x\]

OpenStudy (anonymous):

\[3\sin x=2\cos ^{2}x\] \[\cos ^{2}x+\sin ^{2}x=1\] \[\cos ^{2}x=1-\sin ^{2}x\] \[3\sin x=2(1-\sin ^{2}x)\] \[2\sin ^{2}x+3sinx-2=0\] (sinx+2)(2sinx-1)=0 1)sinx+2=0 => sinx=-2 undefined 2)2sinx-1=0 =>sinx=1/2 =>x=30 or pi/6

OpenStudy (anonymous):

\[-1\le sinx \le1\]